OpenStudy (anonymous):

simplify (3^(1/2)*3^(-1/3))/(2^(-1/2)*2^(1/3))

OpenStudy (saifoo.khan):

\[(3^\frac12 \times3^{\frac-1}{3} \div (2^\frac{-1}{2}\times 2^\frac13)\]Like that?

OpenStudy (anonymous):

yes, except the 2nd exponent listed has a typo

OpenStudy (saifoo.khan):

\[\Huge (3^\frac12 \times 3^{\frac {-1}{3}} \div (2^\frac{-1}{2}\times 2^\frac13)\]

OpenStudy (saifoo.khan):

like that right?

OpenStudy (anonymous):

yes

OpenStudy (saifoo.khan):

\[\huge 3^{\frac13 + \frac{-1}{3}} \div 2^{\frac{-1}{2}+\frac13}\]Solve..

jhonyy9 (jhonyy9):

do you know the property of exponents ?

jhonyy9 (jhonyy9):

when we need multiply the same terms on exponents those exponents are assumed

jhonyy9 (jhonyy9):

1/3 + (-1/3) = ??? is the same like (+1) + (-1) = ???